OpenStudy (anonymous):

Find the derivative of the function by the limit process. f(x) = 7 x

OpenStudy (anonymous):

my guess is it is \[f(x)=\frac{7}{x}\]so \[f'(x)=\lim_{h\to 0}\frac{\frac{7}{x+h}-\frac{7}{x}}{h}\]

OpenStudy (anonymous):

forgetting the limit for the time being, your job is to "simplify" the complex fraction \[\frac{\frac{7}{x+h}-\frac{7}{x}}{h}\]

OpenStudy (anonymous):

ignore the \(h\) in the denominator, we can put that in later subtract \[\frac{7}{x+h}-\frac{7}{x}\] what do you get ?